OpenStudy (anonymous):

2(q + 3) = 4q + q - 9 Show steps please! :)

OpenStudy (nali):

u want to solve for q ?

OpenStudy (anonymous):

I want to figure out what type of equation it is, by solving what I can.

OpenStudy (nali):

ok so what do u need help with

OpenStudy (anonymous):

I got q = 15, but I wanted someone to double check it for me.

OpenStudy (nali):

ok

OpenStudy (nali):

2 (15 + 3) = 4(15) + 15- 9 36= 66 that is not the right answer

OpenStudy (anonymous):

How did you get 36 = 66?

OpenStudy (nali):

by plugging in 15 for every q in the equation and seeing if both sides equal

OpenStudy (anonymous):

Oh..

OpenStudy (nali):

so do u need help?

OpenStudy (anonymous):

15 = 3q?

OpenStudy (nali):

q= 5 good job that is the right answer

OpenStudy (anonymous):

Awesome :) Thanks

OpenStudy (nali):

u r welcome